OpenStudy (anonymous):

Differentiate f(x)=cos^2(x)sin(x) PLEASE SHOW ALL WORK!

OpenStudy (anonymous):

Product rule!!

OpenStudy (anonymous):

chain rule!

OpenStudy (anonymous):

ok chain rule and product rule. you've got to use what you need to get what you want

OpenStudy (anonymous):

use \((fg)'=f'g+g'f\) with \(f(x)=\cos^2(x), f'(x)=2\cos(x)\sin(x), g(x) = \sin(x), g'(x)=\cos(x)\)

OpenStudy (anonymous):

whoa a mistake \[f(x)=\cos^2(x), f'(x)=-2\cos(x)\sin(x)\]

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (anonymous):

i don't know, i didn't do it